OpenStudy (anonymous):

State the reciprocal. -9/8

OpenStudy (lgbasallote):

hint the reciprocal of 2/3 is 3/2 the reciprocal of 4/5 is 5/4 the reciprocal of -2/3 is -3/2 does that help?

OpenStudy (anonymous):

\[-\frac 98 \implies -\frac{1}{\frac 98} \implies \frac 11 \times -\frac 89 \implies?\]

OpenStudy (lgbasallote):

better

OpenStudy (anonymous):

Got it, thanks

OpenStudy (anonymous):

Glad to help... :)

OpenStudy (lgbasallote):

btw..you're deletophilia is getting on my nerves =_= lol

OpenStudy (anonymous):

? Me deleting posts? lol...